I know the order of operations is PEMDAS, but that doesn't seem to help me with this :(
5/28 + 14/15 DIVIDED BY 3/5 MULTIPLIED BY 9/49

Answer by rothauserc(4718)

5/28 + 14/15 DIVIDED BY 3/5 MULTIPLIED BY 9/49 =
(5/28 + 14/15) / (3/5 * 9/49)
LCM for numerator is 28 * 15 = 420
therefore we have,
(75/420 + 392/420) / (27 / 245) = (467 / 420) / (27 / 245)
invert the numerator and denominator of the denominator fraction and multiply it by the numerator fraction
(467 / 420) * (245 / 27) = 10.089506173